### 02 Jan 2024 v2

# Integral Representations of Three Novel Multiple Zeta Functions for Barnes Type: A Probabilistic Approach

Abstract. Integral representation is one of the powerful tools for studying
analytic continuation of the zeta functions. It is known that Hurwitz zeta
function generalizes the famous Riemann zeta function which plays an
important role in analytic number theory. They both have several
multiple versions in the literature.
In this paper, we introduce three novel multiple
zeta functions for Barnes type and study their integral representations
through hyperbolic probability distributions given by Pitman and Yor
(2003, Canad. J. Math., 55, 292β330). The analytically
continued properties of the three multiple zeta functions are also
investigated. Surprisingly, two of them, unlike the previous
results, can extend analytically
to entire functions in the whole complex plane.
Β
Key words and phrases: Integral representation,
Analytical continuation, Barnes zeta function,
Hyperbolic functions, Hurwitz zeta function, Riemann zeta function.
2020
Mathematics Subject Classifications: Primary 11M06, 11M32, 60E05.
Running title: Integral representations of three novel Barnes
multiple zeta
functions
Postal addresses:
Gwo Dong Lin, Institute of Statistical
Science, Academia Sinica, No.β128, Section 2, Academia Road,
Nankang District, Taipei 11529, Taiwan (ROC). (E-mail:
gdlin@stat.sinica.edu.tw)
Chin-Yuan Hu, National Changhua
University of Education, No.β1, Jinde Road, Changhua 50007, Taiwan
(ROC).
(E-mail: buhuua@gmail.com)
Consider the Hurwitz zeta function
where the parameter and the complex variable has the
real part When , reduces to the
famous Riemann zeta function
which plays an important role in analytic number theory (see, e.g.,
Apostal 1976). It is known that each of both zeta functions in (1)
and (2) can extend to a meromorphic function in the whole complex
plane and has a simple pole at where it has residue 1.
As for multidimensional generalizations of (1) and (2), there are
various multiple (-tuple or -fold) zeta functions proposed in
the literature (see, e.g., the survey paper by Matsumoto 2010, the
thesis by Onozuka 2014, and the references therein), where
is an integer. We recall a few of the results in the following:
(i) EulerβZagier multiple (Riemann) zeta function (Euler 1776,
Zagier 1994, Hoffman 1992, Zhao 2000, Matsumoto 2002a, Mehta et al.
2016)
(ii) Multiple Hurwitz zeta function (Kamano 2006, Murty and Sinha
2006, Kelliher and Masri 2008, Mehta and Viswanadham 2017)
or, in general,
(iii) Barnes multiple zeta function (of one variable) (Barnes 1901, 1904, Choi and Quine 1992, Ruijsenaars 2000, Matsumoto 2002a, 2010)
where and are complex parameters satisfying
some suitable conditions (e.g., all of them have positive real
parts).
For further generalizations of the Barnes multiple zeta function
, see Shintani (1977, 1978) and Mellin (1902, 1904) for
the case of one variable, and Matsumoto (2002a, 2002b, 2003a, 2003b,
2005) for the case of several variables. The analytically continued
properties of the above-mentioned multiple zeta functions have been
investigated. For example, the Barnes multiple zeta function
extends to a meromorphic function in and has
simple poles only at (see Barnes 1904, Ruijsenaars
2000 or Matsumoto 2010). The main tools of analytic continuation of
the zeta functions include integral representation, MellinβBarnes
integral formula, EulerβMaclaurin summation, translation formula
(identity), generalized function, etc. For the recent development on
the multiple zeta functions, see the monograph by Zhao (2016).
In this paper, we will introduce the following three variants (a) β
(c) of the Barnes multiple zeta function, mimicking the
Shintani double zeta functions of the form (Shintani 1975, Kim et
al. 2022)
where denotes the number of distinct solutions to the
quadratic congruence equation
Let be an integer and for convenience let
be
positive real numbers. Denote and
As usual, we adopt the
notations and for real
Then, we define
(a) the multiple hyperbolic sinh-moment function
(b) the multiple hyperbolic cosh-moment function
(c) the multiple hyperbolic tanh-moment function
where are integers,
and
are independently and
identically distributed (i.i.d.) uniform random variables on
It is seen that when all the function
in (3) above reduces to the Barnes
multiple zeta function (see Corollary 1 below). We will
provide the integral representations of the above three novel
multiple zeta functions through the hyperbolic probability
distributions given by Pitman and Yor (2003), and investigate their
analytically continued properties including the set of
singularities. Surprisingly, unlike the previous results, the
functions and in (4) and (5) above can extend analytically to
entire functions in The main results are stated in Section
3. Section 2 provides the needed lemmas. Finally, the proofs of the
main results are given in Section 4.
2. Lemmas
We need the following preparatory lemmas. Lemma 1 presents
the basic properties of the gamma function (see, e.g., Apostal 1976,
Chapter 12). Lemma 2 is about Liouvilleβs extension of Dirichlet
integral theorem (see, e.g., Sharma 2005, Chapter 2, or Gibson
1931, p.β492). The crucial Lemma 3 modifies slightly
Cheng and Huβs (2022) Lemma 2.1. The
hyperbolic random variables in Lemma 4 are introduced and
investigated by Pitman and Yor (2003). On the other hand, for the
derivations of the hyperbolic density functions in Lemma 5, see
Cheng and Hu (2022) as well as Pitman and Yor (2003).
Lemma 1 (Gamma function). Let be the
gamma function defined by
Then the function can extend to a meromorphic function
in the whole complex plane and has simple poles at the points
with residue at where
The gamma function satisfies the equation
It has no zeros, and hence the
reciprocal gamma function, , is an entire function in
with zeros at Moreover, if is
not an integer, then
where denotes the greatest integer less than or equal to
Lemma 2 (LiouvilleβDirichlet integral). Let
be an integer and let be a Lebesgue measurable function
on Then we have
provided that the resulting single integral on the right exists.
Here, the parameters and the region of
multiple integral
In particular, if where and
then the following formula holds:
Lemma 3 (Integral transform of a distribution). Let be a random variable with distribution on
denoted and let have finite moments of all orders.
For any nonzero real define the complex-valued function
where Then is an
entire function in
Proof. We sketch the proof below. For each define the
closed square
Denote the integer Then there exists a constant
such that
This implies that the integral transformation is
absolutely convergent and converges uniformly in each compact
subset of Moreover, is a continuous function in
Let be any closed triangle in and recall
that for each fixed real is an entire function in
Then we have, by Cauchyβs theorem,
where is the boundary of Finally,
Moreraβs theorem completes the proof.
Remark 1. Note that for the
function value of becomes a linear combination of moments
of Thus, for simplicity, we might call the integral
transformation a moment function of the distribution
Lemma 4 (Hyperbolic random variables). For
let be three random
variables with characteristic functions
respectively. Then the hyperbolic random variables are infinitely divisible and have finite
moments of all orders.
For
let where are called
hyperbolic -, -, -distributions, respectively.
Lemma 5 (Hyperbolic distributions). For
are absolutely continuous
symmetric distributions about zero and have probability
density functions
respectively, where is the Beta function
In particular, for we have the even density functions:
Throughout the section we assume that is an integer and
that , and are
positive real numbers satisfying Denote
and
Moreover, the
notation ββ means equality in distribution.
Then we have the following results, which as mentioned below
generalize some of the previous results given in Cheng and Hu
(2022).
Theorem 1. Assume that
are independent random variables, where are
hyperbolic sinh random variables defined in Lemma . Then the
multiple hyperbolic sinh-moment function in has an integral representation of the form:
where
Moreover, the following statements hold true.
The zeta function in
can extend
to a meromorphic function in and all the poles are simple.
If the parameter is a positive integer, then the
function in is an analytic
function in except for Precisely,
If the parameter is not an integer, then the zeta
function in has trivial zeros
at and
we have the
following two cases for poles.
For non-negative integer the point is a
simple pole of with residue
For positive integer the point is a
simple pole of with residue
where
If then
and the multiple hyperbolic sinh-moment function in (3) reduces to the Barnes multiple zeta function
Thus, we have the following.
Corollary 1. The Barnes multiple zeta function
has the integral representation
where and are
independent random variables, each
having the density function
Moreover, in extends to a meromorphic function in and has simple poles
at
If, further, and then the multiple
hyperbolic sinh-moment function
reduces to the Hurwitz zeta function Thus, we have the
following.
Corollary 2 (Cheng and Hu 2022). The Hurwitz
zeta function has the integral representation of the
form
where and
has the density function in
Moreover, in extends to a meromorphic function in and has a simple pole
at
If and then
and the multiple hyperbolic sinh-moment function in (3) reduces to
Corollary 3 (Cheng and Hu 2022). The function
in has the integral representation
where and has
the density function
Moreover, in extends to a meromorphic function in and has simple poles
at
Theorem 2. Assume that
are independent random variables, where are
hyperbolic cosh random variables defined in Lemma Then the
multiple hyperbolic cosh-moment function in has an integral representation of the form:
where Moreover, the zeta
function in can extend
analytically to
an entire function in
If then
and the multiple hyperbolic cosh-moment function in (4) reduces to the following zeta function
(compare with the zeta function ). Thus, we have the
following.
Corollary 4. The zeta function
in has the integral representation of the form
where and are
independent random variables, each
having the density function
Moreover, the zeta function in extends to an
entire function in
If, further, and then the zeta
function in (13) reduces to
Thus, we have the following.
Corollary 5 (Cheng and Hu 2022). The zeta
function in has the integral representation
where and has the density function
in Moreover, the zeta function in
extends to an entire function in
If and then
and the multiple hyperbolic cosh-moment function in (4) reduces to
Then we have the following.
Corollary 6 (Cheng and Hu 2022). The zeta
function in has the integral representation
of the form
where and has
the density function
Moreover, the zeta function in extends to an
entire function in
Theorem 3. Assume that
are independent random variables, where
are positive integers and are hyperbolic tanh
random variables defined in Lemma Then the multiple hyperbolic
tanh-moment function in has
an integral representation of the form:
where Moreover, the zeta
function in can extend
analytically to an entire function in
If then and
the multiple hyperbolic tanh-moment function in (5) reduces to
where and
are i.i.d.
uniform random variables on
Thus, we have the
following.
Corollary 7. The multiple hyperbolic
tanh-moment function in
has the integral
representation of the form
where and are i.i.d.
random variables, each having the
density function
Moreover, in extends to an entire function in
If, further, and then the
multiple hyperbolic tanh-moment
function in (21) reduces to
where has the uniform distribution in
Corollary 8 (Cheng and Hu 2022). The multiple
hyperbolic tanh-moment function in
has the integral representation
of the form
where has the density function
in Moreover, in extends to an entire
function in
If and then
and the multiple hyperbolic tanh-moment function
in (21) reduces to
where are i.i.d. uniform random variables on
Corollary 9 (Cheng and Hu 2022). The zeta
function in has the integral
representation
where has the density
function
Moreover, in extends to an entire function
in
Remark 2. In addition to the above single variable cases, it
is also possible to consider the following cases of several
variables, where
and
in which the parameters and
In these cases,
Lemma 2 applies. The results are too complicated to be presented
here.
4. Proofs of the main results
It suffices to prove Theorems 1, 2 and 3, because the corollaries
can be derived easily from the theorems.
Proof of Theorem 1. We first claim that the multiple
series on the RHS of (3) converges absolutely provided To see this, note that by Stirlingβs formula, or more
generally, as
there exist some constants such that
On the other hand, for fixed with take
such that then
Finally, apply the fact that each series
converges.
Multiplying both sides of the
multiple hyperbolic sinh-moment function in (3) by we have
Recall the gamma identity
and the binomial series or, equivalently,
Now, by (29), (30), the above claim and Fubiniβs theorem, we can
rewrite (28) as follows:
Next, consider the product term of the integrand in (31) and write,
using Lemma 4,
where
Inserting (32) in (31) yields, by (29) and Fubiniβs theorem again,
where we apply the symmetric property of the hyperbolic
sinh-distribution. This proves the integral representation in (6) by
dividing both sides of (33) by
Recall that the integral transformation is an entire function in
due to Lemma 3. Therefore, the zeta function in (6) can extend analytically to a meromorphic function
in and all the poles are simple due to Lemma 1. This proves
part (a). Parts (b) and (c) follow from part (a) and Lemma 1 again
immediately. The proof of Theorem 1 is completed.
Proof of Theorem 2. Note first that there is a
relationship between and specifically, we can rewrite
so that the multiple series on the RHS of (4) also converges
absolutely provided Then proceeding along the
same lines in the proof of Theorem 1, we can prove the theorem as
follows. Multiplying both sides of the multiple hyperbolic
sinh-moment function in (4) by
we have
As before, using the gamma identity and the binomial series, we can
rewrite (34) as follows:
Next, consider the product term of the integrand in (35) and write,
using Lemma 4,
where
Inserting (36) in (35) yields, by (29) and Fubiniβs theorem again,
where we apply the symmetric property of the hyperbolic
cosh-distribution. By cancelling the common term in both
sides of (37), we
proves the integral representation in (9).
Recall again that the integral transformation is an entire function in due to
Lemma 3. Therefore, the zeta function in (12) can extend analytically to an entire function in
This completes the proof of Theorem 2.
Proof of Theorem 3. As in the proof of Theorem 2,
the multiple series on the RHS of (5) converges absolutely provided
by noting that
Multiplying both sides of the function in (5) by we have
Then, by (29) and Fubiniβs theorem, we can rewrite the expectation
where we use the fact that for uniform random variable on
Next, recall that
and that
Then it follows from (38)β(41) and the definition of the random
variable that
Cancelling in both sides of (42), we obtain the integral
representation in (20). Finally, Lemma 3 completes the proof of the
theorem.